an organic compound was analysis and found to contain 5.359 gram of carbon , 0.909 gram of hydrogen and three find the empirical and molecular formula if relative mass is 88

1
Expert's answer
2022-01-25T14:45:01-0500

Some information missing.


if the question read:


An organic compound was analysed and found to contain 5.359 gram of carbon , 0.909 gram of hydrogen and 3.636 grams of Oxygen. Find the empirical and molecular formula if relative mass is 88.


Solution


Mole ratio


C:H:O "=\frac {5.359}{12}:\frac {0.909}{1}:\frac{3.636}{16}"


"=0.4466:0.909:0.2273"


Dividing by smallest and rounding to nearest whole number


"C:H:O= 2:4:1"


Empirical formula is "C_2H_4O"

Molar mass =44


Formula mass is twice


Molecular Formula= "C_4H_8O_2"